OpenStudy (anonymous):

determine the slope of a line and the equation of the line in y=mx+b form (2,3) (5,-2) (2,5) (1,7) (-2,-5) (1,7)

OpenStudy (anonymous):

(2,3) (5,-2): y = -4/3x + 19/3 (2,5) (1,7): y = -2x + 9 (-2,-5) (1,7): y = 4x + 3